# The Value of Context: Human versus Black Box Evaluators

### v1

@@ Abstract
Evaluations once solely within the domain of human experts (e.g., medical diagnosis by doctors) can now also be carried out by machine learning algorithms. This raises a new conceptual question: what is the difference between being evaluated by humans and algorithms, and when should an individual prefer one form of evaluation over the other? We propose a theoretical framework that formalizes one key distinction between the two forms of evaluation: Machine learning algorithms are standardized, fixing a common set of covariates by which to assess all individuals, while human evaluators customize which covariates are acquired to each individual. Our framework defines and analyzes the advantage of this customization—the value of context—in environments with very high-dimensional data. We show that unless the agent has precise knowledge about the joint distribution of covariates, the value of more covariates typically exceeds the value of context.
@@ 1 Introduction
In contrast, individuals can often explain their unusual circumstances or characteristics to a human evaluator through conversation. Thus, even if the human evaluator considers fewer inputs than a black box algorithm does, these inputs may be better adapted to the individual being evaluated. Longoni et al. (2019) report that the perception that humans are better able to take into account an individual’s unique situation contributes significantly to patient resistance to AI in healthcare. When, and to what extent, does this difference matter?
@@
In our model, an agent is described by a binary covariate vector and a real-valued type (e.g., the severity of the agent’s medical condition). The type can be written as a function of the covariates, which we henceforth call the type function. Covariates are separated into standard covariates (e.g., medical history, lab tests, imaging scans) and nonstandard covariates (e.g., religious information, genetic data, wearable device data, and financial data). Since in principle there is no limit on the number of nonstandard covariates that can describe a person, our results consider asymptotics as the number of nonstandard covariates grows to infinity.
We suppose that the agent may know how the standard covariates are correlated with the type, but cannot distinguish between the predictive roles of the nonstandard covariates. Formally, the agent has a belief over the type function, and we impose an exchangeability assumption that says that the agent’s prior over these functions is unchanged by permuting the labels and values of the nonstandard covariates. If we interpret the covariates as signals about the agent’s type, then uncertainty about the type function corresponds to uncertainty about the signal structure (à la model uncertainty, e.g., Acemoglu et al. (2015) and Morris and Yildiz (2019)).
@@
First, the black box evaluator observes a larger fraction of the nonstandard covariates than the human evaluator does. Second, the nonstandard covariates observed by the black box evaluator are a pre-specified set of algorithmic inputs, which are fixed across individuals. For example, a designer of a medical algorithm may specify a set of inputs including (among others) blood type, BMI, and smoking status, and train a black box algorithm to learn the mapping from those inputs into the diagnosis. We view the human evaluator as instead uncovering nonstandard covariates during a conversation, where the specific path of questioning may vary across agents. Thus the human evaluator may end up learning about one individual’s sleep schedule and another’s financial situation, where the final set of observed covariates is a function of the agent’s covariate vector.
@@
This comparison essentially reduces to the question of whether the agent prefers an evaluator who observes a larger fraction of (non-targeted) nonstandard covariates about the agent, or an evaluator who observes a smaller but targeted fraction of nonstandard covariates. Towards this comparison, we first introduce a benchmark, which is the expected payoff that the agent would receive if interacting with an evaluator who observes no nonstandard covariates. We define the value of context to be the improvement in the agent’s payoffs under best-case human evaluation, relative to this benchmark. The value of context thus quantifies the extent to which the agent’s payoffs can possibly be improved when the evaluator observes nonstandard covariates suited to that agent.
Our first main result says that under our symmetry assumption on the agent’s prior, the expected value of context vanishes to zero as the number of covariates grows large. Thus even though there may be realizations of the type function given which the value of context is large, in expectation it is not. The contrapositive of this result is that if the expected value of context is high in some application, it must be that our symmetry assumption does not hold, i.e., the agent has some ex-ante knowledge about the predictive roles of the nonstandard covariates.
@@
We next use this result to compare the agent’s expected payoff under human and black box evaluation. We show that when the agent prefers a more accurate evaluation—formally, when the agent’s payoff is convex in the evaluation—the agent should prefer an algorithmic evaluator with access to more covariates over a human evaluator to whom the agent can provide context. And when the agent’s payoff is concave in the evaluation, the conclusion is reversed.
@@
We conclude by examining the role of the symmetry assumption on the agent’s prior, and the extent to which our results depend upon it. First, we study two variations of our main model, in which the symmetry assumption is relaxed: In the first, we suppose that there is a “low-dimensional” set of covariates that relevant for predicting the agent’s type; in the second, we suppose that the agent knows ex-ante the predictive role of certain nonstandard covariates. In both of these settings, our main results extend partially but can also fail: For example, if the set of relevant covariates is sufficiently small that they can be fully disclosed to the evaluator, then the expected value of context typically will not vanish. Finally we show that precise symmetry is not important for our result—it is enough if the informativeness of each individual set of covariates vanishes as the total number of covariates grows large. Together with our main results, these extensions clarify different categories of informational assumptions under which the expected value of context does or does not turn out to be high.
Our model is not meant to be a complete description of the differences between human and black box evaluation. For example, we have abstracted away from human and algorithmic bias (Kleinberg et al., 2017; Gillis et al., 2021), factors such as empathy, and the possibility that the human evaluator has access to information that is not available to the algorithm (e.g., for privacy protection as in Agarwal et al. (2023)). We also suppose that both evaluators form correct conditional expectations, thus abstracting away from the possibility of algorithmic overfitting and of bounded human rationality (e.g., as considered in Spiegler (2020) and Haghtalab et al. (2021)). [^1] The problem of overfitting, while practically important, is a function of how the algorithm is trained. We are interested here in intrinsic differences between the qualitative nature of human and black box evaluation, which are difficult to resolve by training the algorithm differently. We leave extensions of our model that include these other differences to future work.
@@ 1.1 Related Literature
A large literature compares the accuracy of human evaluation with AI evaluation, finding that machine learning algorithms outperform experts in problems including medical diagnosis (Rajpurkar et al., 2017; Jung et al., 2017; Agarwal et al., 2023), prediction of pretrial misconduct (Kleinberg et al., 2017; Angelova et al., 2022), and prediction of worker productivity (Chalfin et al., 2016). These results have led some to predict or call for the replacement of human evaluation with algorithmic evaluation (Obermeyer and Emanuel, 2016). But human evaluation and/or human oversight of algorithmic predictions remains the norm, in part because of user distrust of algorithmic predictions (Jussupow and Heinzl, 2020; Bastani et al., 2022; Lai et al., 2023).
@@
Our formal results are related to the literature on asymptotic learning and agreement across Bayesian agents (Blackwell and Dubins, 1962; Cripps et al., 2008; Acemoglu et al., 2015). Specifically, one can view our main result as bounding (in expectation) the differences in beliefs across Bayesian agents who are given different information. But the asymptotics that we look at are of a different nature from those studied previously. Among other distinctions, we consider asymptotics with respect to a sequence of varying information structures, rather than studying asymptotic beliefs as the total amount of information accumulates.
The agent in our framework has model uncertainty (Acemoglu et al., 2015; Morris and Yildiz, 2019), and the central Assumption 1 constrains the agent’s model uncertainty to take a particular (and new) form motivated by the applications we have in mind. The presence of model uncertainty distinguishes our problem from the related and very interesting work of Di Tillio et al. (2021), which compares the informativeness of an unbiased signal to the informativeness of a “selected” signal whose realization is the maximum realization across i.i.d. unbiased signals. In Di Tillio et al. (2021), the signal structures that are being compared are deterministic and known, while in ours they are random and compared in expectation. [^3] This distinction is important: If, similar to Di Tillio et al. (2021), we modeled covariates as IID signals about the agent’s type, then the expected value of context would not vanish as the number of covariates grew large. See Section 3.2 for more detail.
Finally, our work builds on the literature on persuasion via strategic information disclosure (e.g., Glazer and Rubinstein (2004), Kamenica and Gentzkow (2011)). The model that we study—in which the sender has private information about his type vector, and selectively chooses which elements to disclose to a naive receiver—is closest to models of disclosure of hard information (Dye, 1985; Grossman and Hart, 1980), in particular Milgrom (1981). [^4] A similar model of information is considered in Glazer and Rubinstein (2004) and Antic and Chakraborty (2023). Different from this literature, our sender has uncertainty about how his reports are interpreted, and our focus is not on examining which incentive-compatible reporting strategy is optimal. (Indeed, in our main model we do not require choice of an incentive-compatible reporting strategy, since the receiver updates to the sender’s disclosure as if it were exogenous information. This is primarily for convenience—we show in Section 4.2 that our results extend in a disclosure game.) Our focus is instead on asymptotic limits of belief manipulability as the number of components in the type vector grows large, which is special to our motivation.
@@ 2.1 Setup
We refer to the covariates indexed to $\mathcal{S}=\{1,\dots,s\}$ as standard covariates and the covariates indexed to $\mathcal{N}=\{s+1,\dots,n\}$ as nonstandard covariates. Since in principle there is no limit on the number of covariates that can describe a person, we view the case of infinite covariates as the relevant one, and our results focus on asymptotics as $n$ grows large. [^8] We consider a generalization of our model in Section 5.1 in which a “low-dimensional” subset of these covariates are sufficient for predicting the agent’s type.
@@ 2.2 Evaluation of the agent
There are two evaluators: a black box evaluator, henceforth Black Box (it), and a human evaluator, henceforth Human (she). Both evaluators form an evaluation as an expectation of the agent’s type $y$ given observed covariates, so we will introduce notation for these conditional expectations. For any covariate vector $\mathbb{x}_{n}=(x_{1},\dots,x_{n})$ and subset of nonstandard covariates $A\subseteq\mathcal{N}$, let
@@
Human differs from Black Box in two ways. First, Human has a capacity of $h_{n}=\lfloor\alpha_{h}\cdot n\rfloor$ nonstandard covariates per agent, where $\alpha_{h}<\alpha_{b}$ (i.e., Human cannot process as many inputs as Black Box). Second, Human does not pre-specify which nonstandard covariates to observe, but rather learns these through conversation, and thus potentially observes different nonstandard covariates for each agent. For example, a doctor (evaluator) may pose different questions to different patients (agents) depending on their answers to previous questions. Or a job candidate (agent) might choose to disclose to an interviewer (evaluator) certain nonstandard covariates that put him in a good light.
@@
If however the covariates are disclosed by the agent in a disclosure game, and the evaluator accounts for the strategic nature of this disclosure, then whether (2.3) represents an upper bound will depend on what we assume that the agent knows at the time of disclosure. [^13] This is roughly because the agent can potentially “sneak in” information about the other covariates via the covariates that are revealed. We show in Section 4.2 that if the agent knows his entire covariate vector, then (2.3) need not upper bound every agent’s payoffs. Nevertheless, we present a different quantity that does upper bound the maximum payoff that any agent can obtain in this disclosure game, and show that our main results extend when we replace (2.3) with this quantity. To streamline the exposition we focus on the prior two interpretations (in which the human evaluator either selects the covariates herself or updates to the agent’s disclosures naively), and postpone the discussion of disclosure games to Section 4.2.
@@ 2.3 Value of context
We are interested in settings where the agent does not know the type function, and hence cannot compute the value of context. We give the agent uncertainty about $f$ and characterize the agent’s expected value of context and expected payoffs, integrating over the agent’s belief about $f$ [^15] If we interpret the covariates in our model as signals about the type, then the function relating covariates to type corresponds to the signal structure.
@@
Second, it turns out that the case with uncertainty about $f$ yields a more elegant analysis. Although the value of context can be computed for specific $f$, it typically depends on details of that function, as well as on the agent’s own covariate vector. In contrast, under a condition on the prior belief (defined in the following section) it is possible to draw strong detail-free conclusions.
@@ 2.4 Model Uncertainty
Our leading assumption is that the agent may know how standard covariates impact the type, but has no ex-ante knowledge about the different roles of the nonstandard covariates. Formally, we impose the following.
Definition 2.2 (Finite Exchangeability).
The sequence of random variables $W_{1},W_{2},\dots,W_{n}$ is finitely exchangeable if for every permutation $\pi:\{1,\dots,n\}\rightarrow\{1,\dots,n\}$, the sequences $(W_{1},\dots,W_{n})$ and $(W_{\pi(1)},\dots,W_{\pi(n)})$ have the same joint distribution.
Definition 2.3 (Infinite Exchangeability).
The sequence of random variables $W_{1},W_{2},\dots$ is infinitely exchangeable if for every $n\in\mathbb{N}$ and set of indices $i_{1},\dots,i_{n}$, the sequence $(W_{i_{1}},\dots,W_{i_{n}})$ is finitely exchangeable.
In what follows, for each covariate vector $\mathbb{x}$ we define the random variable $Y_{\mathbb{x}}=f(\mathbb{x})$ (where the randomness is through the type function $f$).
Assumption 1.
Fix any realization of the standard covariates $\mathbb{x}_{\mathcal{S}}=(x_{1},\dots,x_{s})\in\{0,1\}^{s}$. There is an infinitely exchangeable sequence of $[-\overline{y},\overline{y}]$-valued random variables $(\widetilde{Y}_{1},\widetilde{Y}_{2},\dots)$ such that for every $n\in\mathbb{N}$, the sequence
$(Y_{\mathbb{x}_{\mathcal{S}},\mathbb{x}_{-\mathcal{S}}}:\mathbb{x}_{-\mathcal{S}}\in\{0,1\}^{n-s})$
has the same distribution as $(\widetilde{Y}_{1},\dots,\widetilde{Y}_{2^{n-s}})$.
@@
The sequence $(Y_{\mathbb{x}_{\mathcal{S}},\mathbb{x}_{-\mathcal{S}}}:\mathbb{x}_{-\mathcal{S}}\in\{0,1\}^{n-s})$ includes all types associated with the covariate vectors that “complete” $\mathbb{x}_{\mathcal{S}}$ by filling in values for the nonstandard covariates. Assumption 1 says that the joint distribution of these types is ex-ante invariant to permutations of the covariate vectors within the set $\{(\mathbb{x}_{\mathcal{S}},\mathbb{x}_{-\mathcal{S}}):\mathbb{x}_{-\mathcal{S}}\in\{0,1\}^{n-s}\}$. An agent whose prior is given by this joint distribution is thus agnostic about how the labels and values of the nonstandard covariates impact the type.
Besides imposing ex-ante symmetry of the nonstandard covariates, the main content of the assumption is that the unconditional distribution of $y$ in the population of agents is constant across $n$ [^17] This contrasts, for example, with a model in which $x_{1},x_{2},\dots$ were drawn i.i.d. from a type-dependent distribution $F_{y}$. The total quantity of information about $y$ would then be increasing in the number of covariates. In our model it is not. If we interpret each covariate as a bit of information, then under Assumption 1, the number of covariates $n$ can be interpreted as moderating the richness of the informational environment and the potential complexity of the mapping $f$, but not as a measure of the quantity of information. [^18] As we show in Appendix A.1, it is possible to extend our framework so that the type is random conditional on the covariate vector $\mathbb{x}_{n}$; what is crucial is that the size of residual uncertainty does not depend on $n$. In other words, as $n$ grows large, the agent has a more extensive set of words to describe a fixed unknown $y$.
Some simple examples of priors satisfying this assumption are given below. These examples clarify that symmetry of the nonstandard covariates is imposed only on the agent’s prior belief, and not on the realized $f$ [^19] That is, we do not restrict the agent’s prior belief to have support on functions $f$ where $f(x_{1},\dots,x_{n})=f(x_{\pi(1)},\dots,x_{\pi(n)})$ for permutations $\pi:\{1,\dots,n\}\rightarrow\{1,\dots,n\}$.
@@
Let $y\in\{0,1\}$, in which case the space of possible functions $f:\mathcal{X}\rightarrow\mathcal{Y}$ can be identified with $\{0,1\}^{2^{n}}$. Suppose that for each $n$, the agent has a uniform prior on the set of all functions $\{0,1\}^{2^{n}}$. Then Assumption 1 is satisfied.
@@
$\left(f(\mathbb{x}_{\mathcal{S}},\mathbb{x}_{-\mathcal{S}}):\mathbb{x}_{-\mathcal{S}}\in\{0,1\}^{n-s}\right)\sim_{i.i.d.}F.$
Then Assumption 1 is satisfied.
Assumption 1 describes settings in which many different structural relationships between the covariates and the type are plausible (including both ones where the value of context will turn out to be high and low), but ex-ante those relationships are not known. It thus rules out prior knowledge about asymmetries across covariates or covariate values, such as in the following examples.

### v2

@@ Abstract
Machine learning algorithms are now capable of performing evaluations previously conducted by human experts (e.g., medical diagnoses). How should we conceptualize the difference between evaluation by humans and by algorithms, and when should an individual prefer one over the other? We propose a framework to examine one key distinction between the two forms of evaluation: Machine learning algorithms are standardized, fixing a common set of covariates by which to assess all individuals, while human evaluators customize which covariates are acquired to each individual. Our framework defines and analyzes the advantage of this customization—the value of context—in environments with high-dimensional data. We show that unless the agent has precise knowledge about the joint distribution of covariates, the benefit of additional covariates generally outweighs the value of context.
@@ 1 Introduction
In contrast, individuals can often explain their unusual circumstances or characteristics to a human evaluator through conversation. Thus, even if the human evaluator considers fewer inputs than a black box algorithm does, these inputs may be better adapted to the individual being evaluated. The perception that humans are better able to take into account an individual’s unique situation is a significant factor in patient resistance to AI in healthcare (Longoni et al., 2019). Our objective is to understand when, and to what extent, this difference between human and machine evaluation matters.
@@
In our model, an agent is described by a binary covariate vector and a real-valued type (e.g., the severity of the agent’s medical condition). The type can be written as a function of the covariates, which we henceforth call the type function. Covariates are separated into standard covariates (e.g., medical history, lab tests, imaging scans) and nonstandard covariates (e.g., religious information, genetic data, wearable device data, and financial data).
We suppose that the agent may know how the standard covariates are correlated with the type, but cannot distinguish between the predictive roles of the nonstandard covariates. Formally, the agent has a belief over the type function, and we impose two assumptions on the agent’s prior. The first is a symmetry assumption that says that the agent’s prior over these functions is unchanged by permuting the labels and values of the nonstandard covariates. If we interpret the covariates as signals about the agent’s type, then uncertainty about the type function corresponds to uncertainty about the signal structure (à la model uncertainty, e.g., Acemoglu et al. (2015) and Morris and Yildiz (2019)). The second assumption fixes the unpredictability of the agent’s type to be constant in the total number of covariates. We impose this because in many applications, machine learning algorithms have millions of inputs, and yet cannot predict the outcome perfectly. Thus our “many covariates” limit does not represent a situation in which the total of amount of information grows large, but rather one in which the type function can be arbitrarily complex. We view these two assumptions as useful conceptual benchmarks, but subsequently show that neither are essential for our main results (see Section 5 for details).
@@
First, the black box evaluator observes a larger fraction of the nonstandard covariates than the human evaluator does (since humans cannot process as much information as black box algorithms can). Second, the nonstandard covariates observed by the black box evaluator are a pre-specified set of algorithmic inputs, which are fixed across individuals. For example, a designer of a medical algorithm may specify a set of inputs including (among others) blood type, BMI, and smoking status. The black box algorithm learns a mapping from those inputs into the diagnosis. We view the human evaluator as instead uncovering nonstandard covariates during a conversation, where the specific path of questioning may vary across agents. Thus the human evaluator may end up learning about one individual’s sleep schedule but another individual’s financial situation.
@@
This comparison essentially reduces to the question of whether the agent prefers an evaluator who observes a larger fraction of (non-targeted) nonstandard covariates about the agent, or an evaluator who observes a smaller but targeted fraction of nonstandard covariates. [^1] This question is spiritually related to Akbarpour et al. (2024)’s comparison of the network diffusion value of a small number of targeted seeds versus a larger number of randomly selected seeds. Like them, we will find that a larger number of (non-targeted) inputs is superior, but the mechanisms behind these results are very different; in particular, network structure does not play a role in our results. Towards this comparison, we first introduce a benchmark, which is the expected payoff that the agent would receive if interacting with an evaluator who observes no nonstandard covariates. We define the value of context to be the improvement in the agent’s payoffs under best-case human evaluation, relative to this benchmark. The value of context thus quantifies the extent to which the agent’s payoffs can possibly be improved when the evaluator observes nonstandard covariates suited to that agent.
Our first main result says that under our assumptions on the agent’s prior, the expected value of context vanishes to zero as the number of covariates grows large. Thus even though there may be realizations of the type function given which the value of context is large, in expectation it is not. The contrapositive of this result is that if the expected value of context is high in some application, it must be that our assumptions on the prior do not hold, i.e., the agent has some ex-ante knowledge about the predictive roles of the nonstandard covariates.
@@
We next use this result to compare the agent’s expected payoff under human and black box evaluation, when the total number of covariates is sufficiently large. We show that when the agent prefers a more accurate evaluation—formally, when the agent’s payoff is convex in the evaluation—the agent should (eventually) prefer an algorithmic evaluator with access to more covariates over a human evaluator to whom the agent can provide context. And when the agent’s payoff is concave in the evaluation, the conclusion is (eventually) reversed. We view these conclusions as relevant not only in the many-covariates limit: We quantify the number of covariates that is needed for our result to hold, and show that it can be quite small. For example, if the agent’s utility function satisfies mild regularity conditions, the Human evaluator observes 10% of covariates, and the Black Box evaluator observes 90%, of covariates, then our result holds as long as there are at least 14 covariates.
@@
We conclude by examining the role of our assumptions about the agent’s prior, and the extent to which our results depend on them. First, we study two variations of our main model, in which the symmetry assumption is relaxed: In the first, we suppose that there is a “low-dimensional” set of covariates that relevant for predicting the agent’s type; in the second, we suppose that the agent knows ex-ante the predictive role of certain nonstandard covariates. In both of these settings, our main results extend partially but can also fail: For example, if the set of relevant covariates is sufficiently small that they can be fully disclosed to the evaluator, then the expected value of context typically will not vanish. Next we show that our results extend also in a model in which the predictability of the agent’s type is higher in environments with a larger number of covariates (thus relaxing our second asumption on the prior). Finally we provide an abstract learning condition under which our results extend: It is enough for the informativeness of each individual set of covariates to vanish as the total number of covariates grows large. Together with our main results, these extensions clarify different categories of informational assumptions under which the expected value of context does or does not turn out to be high.
Our model is not meant to be a complete description of the differences between human and black box evaluation. For example, we do not consider human or algorithmic bias (Kleinberg et al., 2017; Gillis et al., 2021), explainability (Yang et al., 2024), preferences for empathetic evaluators, or the possibility that the human evaluator has access to information that is not available to the algorithm (e.g., for privacy protection reasons as in Agarwal et al. (2023)). We also suppose that both evaluators form correct conditional expectations, thus abstracting away from the possibility of algorithmic overfitting and of bounded human rationality (e.g., as considered in Spiegler (2020) and Haghtalab et al. (2021)). [^2] The problem of overfitting, while practically important, is a function of how the algorithm is trained. We are interested here in intrinsic differences between the qualitative nature of human and black box evaluation, which are difficult to resolve by training the algorithm differently. We leave extensions of our model that include these other interesting differences to future work.
@@ 1.1 Related Literature
Our paper is situated at the intersection of the literatures on learning (Section 1.1.1) and strategic information disclosure (Section 1.1.2), where our analysis is primarily differentiated from the previous frameworks by our assumption that the agent has model uncertainty (see Section 1.1.1). Our paper is also inspired by a recent empirical literature that compares human and AI evaluation, which we review in Section 1.1.3.
1.1.1 (Asymptotic) Learning
A large literature studies asymptotic learning and agreement across Bayesian agents (Blackwell and Dubins, 1962). Our main result (Theorem 3.1) can be viewed as bounding (in expectation) the differences in beliefs across Bayesian agents who are given different information. As in Vives (1992), Golub and Jackson (2012), Liang and Mu (2019), Harel et al. (2020), and Frick et al. (2023) among others, we quantify the rate of convergence in beliefs. The learning rates that we look at are, however, of a different nature from those studied previously. One important distinction is that these previous papers consider asymptotics as the total amount of information accumulates, while our analysis considers asymptotics with respect to a sequence of information structures that we show are increasingly less informative. A second important difference is that the classic learning models suppose that the agent updates to a signal with a known signal structure, while our agent has uncertainty over the signal structure (as in Acemoglu et al. (2015) and Morris and Yildiz (2019)). Our results characterize the informativeness of this signal in expectation, where the agent’s model uncertainty takes a particular (and new) form motivated by the applications we have in mind.
Finally, our paper is related to Di Tillio et al. (2021), which compares the informativeness of an unbiased signal to the informativeness of a selected signal whose realization is the maximum realization across i.i.d. unbiased signals. Again the key difference is our assumption of model uncertainty—that is, in Di Tillio et al. (2021), the signal structures that are being compared are deterministic and known, while in ours they are random and compared in expectation. In particular, our agent’s prior belief over signal structures can have support on signal processes which are not i.i.d. (for example, it may be that the meaning of one signal is dependent on the meaning of another).
1.1.2 Strategic Information Disclosure
Several literatures study persuasion via strategic information disclosure. Our model—in which the sender has private information about his type vector, and selectively chooses which elements to disclose to a naive receiver—is closest to models of disclosure of hard information (Dye, 1985; Grossman and Hart, 1980), in particular Milgrom (1981). [^3] A similar model of information is considered in Glazer and Rubinstein (2004) and Antic and Chakraborty (2023). The key difference (which follows from our assumption of model uncertainty) is that our sender has uncertainty about how his reports are interpreted. Additionally, our focus is not on examining which incentive-compatible reporting strategy is optimal, [^4] Indeed, in our main model we do not require choice of an incentive-compatible reporting strategy, since the receiver updates to the sender’s disclosure as if it were exogenous information. This is primarily for convenience—we show in Section 4.2 that our results extend in a disclosure game. but instead on asymptotic limits of belief manipulability as the number of components in the type vector grows large. This latter focus is special to our motivating applications.
1.1.3 Human vs AI Evaluation
Recent empirical papers compare the accuracy of human evaluation with AI evaluation, finding that machine learning algorithms outperform experts in problems including medical diagnosis (Rajpurkar et al., 2017; Jung et al., 2017; Agarwal et al., 2023), prediction of pretrial misconduct (Kleinberg et al., 2017; Angelova et al., 2022), and prediction of worker productivity (Chalfin et al., 2016). Nonetheless, many individuals continue to distrust algorithmic predictions (Jussupow and Heinzl, 2020; Bastani et al., 2022; Lai et al., 2023). These findings motivate our goal of understanding whether individuals should prefer human evaluators, and when instead the replacement of human evaluation with algorithmic evaluation is welfare-improving for users, as suggested in Obermeyer and Emanuel (2016) among others.
@@ 2.1 Setup
We refer to the covariates indexed to $\mathcal{S}=\{1,\dots,s\}$ as standard covariates and the covariates indexed to $\mathcal{N}=\{s+1,\dots,n\}$ as nonstandard covariates.
@@ 2.2 Evaluation of the agent
There are two evaluators: a black box evaluator, henceforth Black Box (it), and a human evaluator, henceforth Human (she). Both form evaluations as an expectation of the agent’s type $y$ given observed covariates, so we will introduce notation for these conditional expectations. For any covariate vector $\mathbb{x}_{n}=(x_{1},\dots,x_{n})$ and subset of nonstandard covariates $A\subseteq\mathcal{N}$, let
@@
Human differs from Black Box in two ways. First, Human has a capacity of $h_{n}=\lfloor\alpha_{h}\cdot n\rfloor$ nonstandard covariates per agent, where $\alpha_{h}<\alpha_{b}$ (i.e., Human cannot process as many inputs as Black Box). Second, Human does not pre-specify which nonstandard covariates to observe, but rather learns these through conversation, and thus potentially observes different nonstandard covariates for each agent. For example, a doctor (evaluator) may pose different questions to different patients (agents) depending on their answers to previous questions. Or a job candidate (agent) might choose to disclose to an interviewer (evaluator) certain nonstandard covariates that are favorable to him.
@@
If however the covariates are disclosed by the agent in a disclosure game, and the evaluator accounts for the strategic nature of this disclosure, then whether (2.3) represents an upper bound will depend on what we assume that the agent knows at the time of disclosure. [^13] This is roughly because the agent can potentially “sneak in” information about the other covariates via the covariates that are revealed. We show in Section 4.2 that if the agent knows his entire covariate vector, then (2.3) need not upper bound every agent’s payoffs. Nevertheless, we present a different quantity that does upper bound the maximum payoff that any agent can obtain in this disclosure game, and show that our main results extend when we replace (2.3) with this quantity. To streamline the exposition we focus on the prior two interpretations (in which the human evaluator either selects the covariates herself or updates to the agent’s disclosures naively), and discuss disclosure games in Section 4.2.
@@ 2.3 Value of context
In what follows, we give the agent uncertainty about $f$ and characterize the agent’s expected value of context and expected payoffs, integrating over the agent’s belief about $f$ [^15] If we interpret the covariates in our model as signals about the type, then the function relating covariates to type corresponds to the signal structure.
@@
Second, the case with uncertainty about $f$ turns out to yield a more elegant analysis than the one in which $f$ is known. That is, although the value of context for specific functions $f$ depends on details of that function and on the agent’s own covariate vector, there is a large class of prior beliefs (described in the following section) for which it is possible to draw strong detail-free conclusions about the expected value of context.
@@ 2.4 Model Uncertainty
We impose two assumptions on the agent’s prior belief about $f$. Together, these assumptions deliver a setting in which many different structural relationships between the covariates and the type are possible (including both ones where the value of context is high and low), but ex-ante those relationships are not known.
@@
The first assumption says that while the agent may know how standard covariates impact the type, he has no ex-ante knowledge about the roles of the nonstandard covariates.
Assumption 1 (Exchangeability).
For every realization of the standard covariates $(x_{1},\dots,x_{s})$, the sequence
$(Y^{n}_{1},\dots,Y^{n}_{2^{n-s}})\equiv(f(x_{1},\dots,x_{s},x_{s+1},\dots,x_{n}):(x_{s+1},\dots,x_{n})\in\{0,1\}^{n-s})$ (2.4)
is finitely exchangeable.
@@
The set $\{(f(x_{1},\dots,x_{s},x_{s+1},\dots,x_{n}):(x_{s+1},\dots,x_{n})\in\{0,1\}^{n-s}\}$ ranges over all covariate vectors that share the standard covariate values $(x_{1},\dots,x_{s})$. Assumption 1 says that the joint distribution of these types is ex-ante invariant to permutations of the labels and values of the nonstandard covariates. An agent whose prior satisfies Assumption 1 is thus agnostic about how the nonstandard covariates impact the type.
While our assumption of no prior knowledge about the role of nonstandard covariates is strong, it is consistent with our interpretation of the nonstandard covariates as those covariates for which there is little historical data about correlations. For example, if it were known that a higher GPA positively correlates with on-the-job performance, but not how a large number of social media followers predicts performance, then we would think of GPA as a standard covariate and the number of social media followers as a nonstandard covariate.
Assumption 1 does not restrict how the agent’s prior varies with $n$, the number of covariates. In a model in which $x_{1},x_{2},\dots$ were drawn i.i.d. from a type-dependent distribution $F_{y}$, the total quantity of information about $y$ would increase in the number of covariates, and the evaluator’s uncertainty about $y$ would vanish as $n$ grew large. This does not seem descriptive of real applications: credit scoring algorithms and healthcare algorithms use millions of covariates, but there remains substantial residual uncertainty about the agent’s type. We take the opposite extreme in which the predictability of the agent’s type is a primitive of the setting, which is held constant for all $n$. In our model, $n$ is not a measure of the total quantity of information, but instead moderates the richness of the informational environment and the potential complexity of the mapping $f$. Loosely speaking, as $n$ grows large, the agent has a more extensive set of words to describe a fixed unknown $y$ [^17] As $n$ grows large, the smallest possible informational size of each covariate (in the sense of McLean and Postlewaite (2002)) vanishes. But we do not require each covariate to be equally informationally relevant in the realized function. So, for example, $f(x_{1},\dots,x_{n})=x_{1}$ can be in the support of the agent’s beliefs for $n$ arbitrarily large (see Example 7).
Assumption 2 (Constant Unpredictability of $Y$).
Fix any realization of the standard covariates $(x_{1},\dots,x_{s})$, and let $(Y^{n}_{1},\dots,Y_{2^{n-s}}^{n})$ be defined as in (2.4) for each $n\in\mathbb{Z}_{+}$. Then for every pair $n^{\prime}>\tilde{n}$, the sequence $(Y^{\tilde{n}}_{1},\dots,Y^{\tilde{n}}_{2^{\tilde{n}-s}})$ and the truncated sequence $(Y^{n^{\prime}}_{1},\dots,Y^{n^{\prime}}_{2^{\tilde{n}-s}})$ are identically distributed.
The statement of Assumption 2 formally depends on the ordering of types within the vector $(Y^{n^{\prime}}_{1},\dots,Y^{n^{\prime}}_{2^{n^{\prime}}})$, since this determines which types appear in the truncated sequence $(Y^{n^{\prime}}_{1},\dots,Y^{n^{\prime}}_{2^{n}})$. But if we further impose Assumption 1 (and we will always impose these assumptions jointly), then the ordering of types is irrelevant: That is, when Assumption 2 holds for one such ordering, it will hold for all orderings.
It is important to note that in our model, Assumptions 1 and 2 are placed ex-ante on the agent’s prior, and not ex-post on the realized function $f$. For example, the function $f(x_{1},\dots,x_{n})=x_{1}$, which says that the only covariate that matters is $x_{1}$, is strongly asymmetric ($x_{1}$ is differentiated from the other covariates) and also features a single “large” covariate (the realization of $x_{1}$ completely determines $y$). Our assumptions do not rule out the possibility of this function. Rather, they require that if this function is considered possible, then certain other functions are as well. [^18] Assumption 1 requires that for every permutation $\pi:\{0,1\}^{n}\rightarrow\{0,1\}^{n}$, the function $g_{\pi}$ satisfying $g_{\pi}(x_{1},\dots,x_{n})=f(\pi(x_{1},\dots,x_{n}))$ is also in the support of the agent’s beliefs.
Simple examples of priors satisfying these two assumptions are given below.
@@
Let $y\in\{0,1\}$, in which case the space of possible functions $f:\mathcal{X}\rightarrow\mathcal{Y}$ can be identified with $\{0,1\}^{2^{n}}$. Suppose that for each $n$, the agent has a uniform prior on the set of all functions $\{0,1\}^{2^{n}}$. Then Assumptions 1 and 2 are satisfied.
@@
$\left(f(x_{1},\dots,x_{s},x_{s+1},\dots,x_{n}):(x_{s+1},\dots,x_{n})\in\{0,1\}^{n-s}\right)\sim_{i.i.d.}F.$
Then Assumptions 1 and 2 are satisfied.
Priors that violate these assumptions include the following.